OpenStudy (anonymous):

Joanne buys one kilogram of French Roast and one kilogram of Boxer Blend from the Java Hut. Jacob buys two kilogram of French Roast and three kilograms of Boxer Blend. If Joanne spends $14 for her purchase, and Jacob spends $36 for his, what is the price of one kilogram of Boxer Blend coffee?

OpenStudy (texaschic101):

f = french roast b = boxer blend f + b = 14 --> f = 14 - b 2f + 3b = 36 I got you started...now just sub 14 - b in for f in the 2nd equation and you will find the cost of boxer blend (b). Easy enough.